A 0.5-L unbuffered solution needs the pH adjusted from 4 to 3.5. How many microliters of a 2 molar HCl solution need to be added to adjust the pH?

Report your answer to four decimal places.
Answer #1

We know that concentration of H+ = 10-pH

Initial concentration of H+ = 10-4 M

Final concentration of H+ = 10-3.5 = 3.16 * 10-4 M

Additional concentration of H+ = final - initial = 3.16 * 10-4 - 1 * 10-4 = 2.16 * 10-4 M

moles of H+ needed = Volume * molarity = 0.5 * 2.16 * 10-4 = 1.08 * 10-4 moles

Volume of HCl * Molarity of HCl = moles of H+

V * 2 = 1.08 * 10-4

SOlving, V = 5.4 * 10-5 Liters = 54 microliters